directions: write an equation for the ellipse with each set of characteristics
Problem: vertices (4,3) (4,-9); length of minor axis-8

Answer by lwsshak3(11628)

directions: write an equation for the ellipse with each set of characteristics
Problem: vertices (4,3) (4,-9); length of minor axis-8
**
Given data shows ellipse has a vertical major axis with equation of standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1, a>b, with (h,k) being the (x,y) coordinates of the center.
..
solving for a, b, and coordinates of the center:
..
x-coordinate of center=4
y-coordinate of center=(3-9)/2=-6/2=-3
center: (4,-3)
..
length of major axis=3-(-9)=12=2a
a=6
a^2=36
..
length of minor axis=8=2b
b=4
b^2=16
..
Equation of given ellipse: (x-4)^2/16+(y+3)^2/36=1